# SegmentDreamer: Towards High-fidelity Text-to-3D Synthesis with Segmented Consistency Trajectory Distillation

### v1

@@ 4.2 Reframing SDS through SCTD
We address the previous limitations by reframing SDS through the lens of SCTD. Specifically, we first define our consistency function as follows: as existing segmented consistency models [44, 33] do, we partition the entire time-step range $[0,T]$ into $N_{s}$ sub-intervals (i.e., $N_{s}+1$ edge points), represented by $\{[s_{m},s_{m+1})|m\in\{0,\cdots,N_{s}-1\},s_{0}=0,s_{N_{s}}=T\}$ (the segmentation strategy will be discussed Sec. 4.3). We define our consistency function $\boldsymbol{G_{\theta}}$ using Eq. 5, aiming to enforce consistency within each sub-trajectory. i.e., $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y})=\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{s},s,s_{m},\mathbf{y})$ for $s_{m}\leq s<t\leq s_{m+1}$. For brevity, we abbreviate $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y})$ as $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\mathbf{z}_{t},t,\mathbf{y})$. Notably, our consistency function $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ avoids the issues inherent in GCS, as analyzed in Sec. 4.1 and Supp. B, since at each training stage, the target time step $s_{m}$ is always uniquely determined.
@@
where $b(t)=\frac{\omega(t)}{(\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda)^{2}}$ and $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{t},t,s,\mathbf{y})$. The derivation is given in App. C. The first term on the r.h.s. of Eq. 8 enforces self-consistency across the conditional denoising sub-trajectory, referred to as the “self-consistency constraint”. The second term enforces cross-consistency between the unconditional and conditional denoising sub-trajectories, termed the “cross-consistency constraint”. The third term adheres to the naming conventions in CSD [53].
As observed, Eq. 8 explicitly defines the relationship between the self-consistency constraint and the cross-consistency constraint, effectively addressing the imbalance problem present in previous CD-based methods [48, 22]. Furthermore, as demonstrated in Sec. 4.3, SCTD provides a tighter upper bound on distillation error, enhancing the upper limit of our algorithm.
@@ 4.3 SCTD Sampling
Sampling Design. A natural question arises: according to Eq. 8, does minimizing $\mathcal{L}_{SDS}$ necessarily enforce both self- and cross-consistency? The answer is no. This can be easily understood by noting that one cannot necessarily achieve $x_{1}=x_{2}$ and $y_{1}=y_{2}$ when $||x_{1}-x_{2}+\omega(y_{1}-y_{2})||_{2}^{2}=0$. While Eq. 8 provides valuable insights, applying it to text-to-3D tasks remains challenging. To address this, we introduce the following modifications. First, we observe that the condition $\mathbf{y}$ appears in all three loss terms in Eq. 8, which increases computational cost particularly when using Perp-Neg [1] or ControlNet [56]. To mitigate this, we apply an equivalent transformation to it:
@@
The derivation is given in Supp. D.
We follow the suggestion of CSD [53], omitting the generative prior from Eq. 9. To enforce the self- and cross-consistency, we consider a more strict constraint on Eq. 8, defining our SCTD sampling as follows:
@@
where $\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}$ is obtained through multi-step deterministic sampling from $\mathbf{z}_{0}$, which is discussed in Sec. 4.4. $\mathbf{z}_{s_{m}}=\alpha_{s_{m}}\mathbf{z}_{0}+\sigma_{s_{m}}\boldsymbol{\epsilon}^{*}$ with $\boldsymbol{\epsilon}^{*}$ fixed for the entire training. $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s,\mathbf{y})$ where $\boldsymbol{\Phi}$ is implemented by a first-order ODE solver in this work. $sg(\cdot)$ represents stopping gradient backpropagation.
@@
9:     Obtain $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)$,$\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ with Eq. 5
10:     Obtain $\nabla_{\theta}\mathcal{L}_{\text{SCTD}}$ with Eq. 10 and update $\boldsymbol{\theta}$.
11:end while
Trajectory Segmentation Strategy. We propose two heuristic trajectory segmentation strategies: equal segmentation and monotonically increasing segmentation. In the first strategy, the entire time-step range $[0,T]$ is uniformly divided into $N_{s}$ segments, each with an equal duration, i.e., $s_{m+1}-s_{m}=\frac{T}{N_{s}}$. In the second strategy, we adopt a monotonically increasing interval segmentation method, where the $N_{s}$ intervals satisfy the condition: $s_{m+1}-s_{m}=t_{\tau}+\frac{2m(T-N_{s}t_{\tau})}{N_{s}(N_{s}-1)},m\in\{0,\cdots,N_{s}-1\}$. A comparison of these two strategies is provided in Sec. 6.
@@ 4.4 Swift and Stable Optimization Pipeline
This strategy enable us to make a trade-off between generation quality and optimization speed, as evidenced in Sec. 6.
Consistency function Approximation. For faster optimization, we propose approximating $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)$ as $\mathbf{z}_{s_{m}}$, leveraging the theoretical invertibility of the unconditional PF-ODE trajectory defined in Eq. 11. However, since both $\boldsymbol{\Phi}$ and $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ are both implemented in a first-order ODE solver in this paper, perfect invertibility cannot be guaranteed. Nevertheless, we find that this approximation strategy not only eliminates the need to compute the U-Net Jacobian—reducing computational cost—but also improves optimization stability, as demonstrated in Sec. 6.
@@ 4.5 Theoretical Analysis of Distillation Error
We theoretically prove that our SCTD loss leads to a tighter upper bound on the distillation error, explaining why SCTD achieves superior generative results. Specifically, the SDS loss can be reformulated as:
@@
Based on Eq. 12, we have the following theorem hold:
@@
Given a sub-trajectory $[s_{m},s_{m+1}]$, let $\Delta t=max_{t,s\in(s_{m},s_{m+1}]}\{|t-s|\}$. We assume $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}$ satisfies the Lipschitz condition. If there exists a $\mathbf{z}_{0}$ satisfying $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})=\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$ for $\forall t,s\in[s_{m},s_{m+1}]$ and a true image $\mathbf{z}^{data}\sim\mathbb{P}_{data}(\mathbf{z})$ corresponding to $\mathbf{z}_{0}$, we have
$\sup_{t,s\in[s_{m},s_{m+1})}||\mathbf{z}_{0}-\mathbf{z}^{data}||\\ =\mathcal{O}(\Delta t)(s_{m+1}-s_{m}),$ (13)
@@
The full proof, inspired by consistency model-related works [27, 44, 48, 22], is provided in App. E. ∎
Theorem 1 shows that our SCTD loss achieves an arbitrarily small distillation error when the ODE solver’s step size $\Delta_{t}$ and the segment length $s_{m+1}-s_{m}$ are sufficiently small. By choosing an appropriate segment number $N_{s}$, we obtain a tighter error upper bound, improving upon the bound $\mathcal{O}(\Delta t)(T)$ in CDS [48] and $\mathcal{O}(\Delta t)(T-e)$ in GCS [22]. This refinement enables more accurate gradient for optimizing the 3D representation $\boldsymbol{\theta}$, leading to higher-quality generative results.
@@ 5 Experiments
Figure 4: Qualitative comparisons with two SD-based metheds: DreamFusion [31] and LucidDreamer [23], and two CD-based methods: Consistent3D [48] and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields high-quality results within the shortest time. Please zoom in for details.
@@ 5.1 Implementation Details
Our text-to-3D algorithm is implemented in PyTorch and trained on an A100 GPU for 5,000 iterations using the Adam optimizer [19]. We utilize Stable Diffusion 2.1 [34] as our foundational generative model and employ Point-E [30] to initialize 3D Gaussians. Hyperparameters related to 3D Gaussian Splatting (3DGS), such as learning rates and camera poses, are set following the guidelines in previous works [23, 22]. $t\sim\mathcal{U}(20,500+t_{warm})$ where $t_{warm}$ decreases linearly from 480 to 0 during the first 1,500 epochs.
@@ 5.2 Text-to-3D Generation
We also compare our SCTD loss with existing CD-based losses, CDS [48] and GCS [22]. For a fair comparison, we use Stable Diffusion 2.1 as the generative model and keep the random seed, guidance scale, and 3DGS-related parameters consistent across all methods. As shown in Fig. 5, CDS struggles to provide effective conditional guidance at a normal CFG scale, as also noted in [22]. Even at a higher CFG scale, its results remain unsatisfactory due to the lack of cross-consistency in its distillation loss. GCS, meanwhile, suffers from overexposure and artifacts due to excessive conditional guidance. While it attempts to address this with Brightness-Equalized Generation (BEG), artifacts persist (see Fig. 5, zoom in for details). In contrast, our SCTD mitigates ineffective conditional generation, overexposure, and artifacts, striking a better balance between self-consistency and cross-consistency.
Figure 5: Visual comparisons with CDS [48], GCS [22], and GCS+BEG [22]. Artifacts are marked with red circles.
Figure 6: Ablation study of the trajectory segmentation strategy with segment numer $N_{s}$ and time theshold $t_{\tau}$. Prompt: a DSLR photo of the Imperial State Crown of England.
@@
Quantitative Comparison. We assess text alignment and human preference for each text-to-3D method in Tab. 1, where we generate 3D models from 40 prompts selected from the DreamFusion Gallery  [^1] https://dreamfusion3d.github.io/gallery.html and compute the average CLIP Score and ImageReward (IR) score across 120 camera views per model. Following Connect3D, we compute the FID score between 40 objects (each renders 120 views) and 50,000 images generated by Stable Diffusion 2.1. As shown, our method achieves the best text alignment, human preference, and fidelity while requiring the least time, showing its superior performance. The 40 prompts are given in the supplementary material.
@@
DreamFusion[31] 28.47 -0.004 60 140.84 4.73 4.87 4.90
LucidDreamer[23] 29.99 0.006 45 121.80 2.93 2.98 2.93
Consistent3D[48] 30.60 0.004 140 113.61 4.14 3.88 3.93
ConnectCD [22] 30.73 0.018 80 112.61 1.63 1.80 1.93
Ours 30.88 0.020 38 110.45 1.57 1.47 1.33
User Study. We conduct a user study with 40 volunteers to assess the generation quality for each algorithm. Each volunteer reviews eight groups of five 360-degree rendered videos and ranks them from the perspective of text alignment (Q1), object realism (Q2), and object details (Q3, e.g., artifact, brightness, Janus problem). To minimize bias, video order is randomized. As shown in Tab. 1, users consistently favor our method across all criteria.
6 Ablation Study
Effect of Trajectory Segmentation Strategy We visualize the effect of different trajectory segmentation strategies applied to our algorithm in Fig. 6. When the segment number $N_{s}$ and the time threshold $t_{\tau}$ are fixed, we observe that the segmentation strategy has minimal impact on the generated results. This is primarily because the segmentation strategy essentially determines $\sup|t-e|$. However, the timesteps $t$ and $s$ are randomly selected, $|t-s_{m}|$, $|t-s|$, and $|s-s_{m}|$ vary randomly during each training iteration, with their values potentially being large or small.
Effect of Segment Number $N_{s}$. We visualize the effect of different segment numbers $N_{s}$ in Fig. 6. As observed, a small $N_{s}$, regardless of the segmentation strategy, consistently results in overly smoothed outputs. As $N_{s}$ increases, the generation quality improves. However, when $N_{s}$ becomes too large, e.g. 10, while more details emerge, the overall model representation becomes vague, which is also mentioned in LucidDreamer [23]. We find that $N_{s}=5$ performs well in most cases.
Effect of time threshold $t_{\tau}$. As shown in Fig. 6, a large $t_{\tau}$ also leads to overly smoothed results for both segmentation strategies. This may be because a limited number of deterministic sampling steps in the diffusion process fails to effectively preserve the information of the initiation point $\mathbf{z}_{0}$, highlighting the necessity of incorporating multi-step deterministic sampling.
Validation of Approximation Strategy. We analyze the impact of replacing $\boldsymbol{G}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ with $\mathbf{z}_{s_{m}}$ in Fig. 7. As shown, $\boldsymbol{G}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ introduces slight overexposure and geometry distortion in generation results. In contrast, $\mathbf{z}_{s_{m}}$ not only produces better generation results but also reduces the training time. This suggests a gap between theory and practice, resembling the omission of U-Net Jacobian in DreamFusion [31].

### v2

@@ 4.2 Reframing SDS through SCTD
We address the previous limitations by reframing SDS through the lens of SCTD. Specifically, we first define our consistency function as follows: as existing segmented consistency models [44, 33] do, we partition the entire time-step range $[0,T]$ into $N_{s}$ sub-intervals (i.e., $N_{s}+1$ edge points), represented by $\{[s_{m},s_{m+1})|m\in\{0,\cdots,N_{s}-1\},s_{0}=0,s_{N_{s}}=T\}$ (the segmentation strategy is discussed in Sec. 4.3). We define our consistency function $\boldsymbol{G_{\theta}}$ using Eq. 5, aiming to enforce consistency within each sub-trajectory, i.e., $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y})=\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{s},s,s_{m},\mathbf{y})$ for $s_{m}\leq s<t\leq s_{m+1}$. For brevity, we abbreviate $\boldsymbol{G}_{\boldsymbol{\theta}}(\mathbf{z}_{t},t,s_{m},\mathbf{y})$ as $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}(\mathbf{z}_{t},t,\mathbf{y})$. Notably, our consistency function $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ avoids the issues inherent in GCS, as analyzed in Sec. 4.1 and Supp. B, because at each training stage, the target time step $s_{m}$ is always uniquely determined.
@@
where $b(t)=\frac{\omega(t)}{(\alpha_{s_{m}}\int_{\lambda_{t}}^{\lambda_{s_{m}}}e^{-\lambda}\mathrm{d}\lambda)^{2}}$ and $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\mathbf{z}_{t},t,s,\mathbf{y})$. The derivation is given in App. D. The first term on the r.h.s. of Eq. 8 enforces self-consistency across the conditional denoising sub-trajectory, referred to as the ”self-consistency constraint.” The second term enforces cross-consistency between the unconditional and conditional denoising sub-trajectories, termed the ”cross-consistency constraint.” The third term adheres to the naming conventions in CSD [53]. As observed, Eq. 8 explicitly defines the relationship between the self-consistency constraint and the cross-consistency constraint. Furthermore, as demonstrated in Sec. 4.3, SCTD provides a tighter upper bound on the distillation error.
@@ 4.3 SCTD Sampling
Sampling Design. A natural question arises: according to Eq. 8, does minimizing $\mathcal{L}_{SDS}$ necessarily enforce both self- and cross-consistency? The answer is no. This is because one cannot necessarily achieve $x_{1}=x_{2}$ and $y_{1}=y_{2}$ when $||x_{1}-x_{2}+\omega(y_{1}-y_{2})||_{2}^{2}=0$. Additionally, we observe that the condition $\mathbf{y}$ appears in all three loss terms in Eq. 8, which increases the computational cost, particularly when using Perp-Neg [1] or ControlNet [56].
To improve the computational efficiency, we perform the following equivalent transformation on Eq. 8:
@@
The derivation is given in Supp. E.
To enforce self- and cross-consistency, we follow the suggestion from CSD [53] to omit the generative prior from Eq. 9, and then impose a stricter constraint on Eq. 8:
@@
where $\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}}$ is obtained through multi-step deterministic sampling from $\mathbf{z}_{0}$, as discussed in Sec. 4.4. $\mathbf{z}_{s_{m}}=\alpha_{s_{m}}\mathbf{z}_{0}+\sigma_{s_{m}}\boldsymbol{\epsilon}^{*}$, with $\boldsymbol{\epsilon}^{*}$ fixed throughout the entire training. $\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}}=\boldsymbol{\Phi}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,s,\mathbf{y})$, where $\boldsymbol{\Phi}$ is implemented by a first-order ODE solver in this work. $sg(\cdot)$ represents the stopping of gradient backpropagation.
@@
9:  Obtain $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\emptyset)$ and $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ with Eq. 5
10:  Obtain $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{SCTD}}$ with Eq. 10 and update $\boldsymbol{\theta}$.
Trajectory Segmentation Strategy. We propose two heuristic trajectory segmentation strategies: equal segmentation and monotonically increasing segmentation. In the first strategy, the entire time-step range $[0,T]$ is uniformly divided into $N_{s}$ segments, i.e., $s_{m+1}-s_{m}=\frac{T}{N_{s}}$. In the second strategy, we adopt a monotonically increasing interval segmentation method, where the $N_{s}$ intervals satisfy that $s_{m+1}-s_{m}=t_{\tau}+\frac{2m(T-N_{s}t_{\tau})}{N_{s}(N_{s}-1)}$ for $\in\{0,\cdots,N_{s}-1\}$. A comparison of the two strategies is presented in Tab. 2.
@@ 4.4 Swift and Stable Optimization Pipeline
This strategy enables us to make a trade-off between generation quality and optimization speed, as evidenced in Tab. 2.
Consistency function Approximation. For faster optimization, we propose approximating $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\emptyset)$ as $\mathbf{z}_{s_{m}}$, leveraging the theoretical invertibility of the unconditional PF-ODE trajectory defined in Eq. 11. However, since both $\boldsymbol{\Phi}$ and $\boldsymbol{G}_{\boldsymbol{\theta}}^{m}$ are implemented using a first-order ODE solver in this paper, perfect invertibility cannot be guaranteed. Nevertheless, we find that this approximation strategy not only eliminates the need to compute the U-Net Jacobian—reducing computational cost—but also improves optimization stability, as demonstrated in Tab. 2.
@@ 4.5 Theoretical Analysis of Distillation Error
We theoretically prove that our SCTD loss leads to a tighter upper bound on the distillation error, which explains why SCTD achieves superior generative results. Specifically, the SDS loss can be reformulated as:
@@
Based on Eq. 12, we have the following theorem:
@@
Given a sub-trajectory $[s_{m},s_{m+1}]$, let $\Delta t=max_{t,s\in(s_{m},s_{m+1}]}\{|t-s|\}$. We assume $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}$ satisfies the Lipschitz condition. If there exists a $\mathbf{z}_{0}$ satisfying $\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})=\boldsymbol{G}^{m}_{\boldsymbol{\theta}}(\hat{\mathbf{z}}_{s}^{\boldsymbol{\Phi}},s,\mathbf{y})$ for $\forall t,s\in[s_{m},s_{m+1}]$ and a true image $\mathbf{z}^{data}\sim\mathbb{P}_{data}(\mathbf{z})$ (corresponding to $\mathbf{z}_{0}$), we have
@@
The full proof, inspired by consistency model-related works [27, 44, 48, 22], is provided in App. F. ∎
Theorem 1 shows that our SCTD loss achieves an arbitrarily small distillation error when the ODE solver’s step size $\Delta_{t}$ and the segment length $s_{m+1}-s_{m}$ are sufficiently small. By choosing an appropriate segment number $N_{s}$, we obtain a tighter error upper bound, which improves upon the bound $\mathcal{O}(\Delta t)(T)$ in CDS [48] and $\mathcal{O}(\Delta t)(T-e)$ in GCS [22]. This refinement enables more accurate gradients for optimizing the 3D representation $\boldsymbol{\theta}$, leading to higher-quality generative results.
@@ 5 Experiments
Figure 4: Qualitative comparison results of two SD-based metheds: DreamFusion [31] and LucidDreamer [23], and two CD-based methods: Consistent3D [48] and ConnectCD [22]. CFG scales are set to 100, 7.5, 20$\sim$40, 7.5, 7.5, respectively. Our approach yields high-quality results within the shortest time. Please zoom in for details.
@@ 5.1 Implementation Details
Our algorithm is implemented in PyTorch and trained on an A100 GPU for 5,000 iterations using the Adam optimizer [19]. We utilize Stable Diffusion 2.1 [34] as our foundational generative model and employ Point-E [30] to initialize 3D Gaussians. Hyperparameters related to 3D Gaussian Splatting (3DGS), such as learning rates and camera poses, are set following the guidelines in previous works [23, 22]. We have $t\sim\mathcal{U}(20,500+t_{warm})$, where $t_{warm}$ decreases linearly from 480 to 0 during the first 1,500 epochs.
@@ 5.2 Text-to-3D Generation
We also compare our SCTD loss with existing CD-based losses, namely CDS [48] and GCS [22]. For a fair comparison, we use Stable Diffusion 2.1 as the generative model and keep the random seed, guidance scale, and 3DGS-related parameters consistent across all methods. As shown in Fig. 5, CDS struggles to provide effective conditional guidance at a normal CFG scale, as also noted in [22]. Even at a higher CFG scale, its results remain unsatisfactory due to the lack of cross-consistency in its distillation loss. GCS, meanwhile, suffers from overexposure and artifacts due to excessive conditional guidance. While it attempts to address this with Brightness-Equalized Generation (BEG), artifacts persist (see Fig. 5; zoom in for details). In contrast, our SCTD mitigates ineffective conditional generation, overexposure, and artifacts, striking a better balance between self-consistency and cross-consistency.
Figure 5: Visual comparison results of CDS [48], GCS [22], and GCS+BEG [22]. Artifacts are marked with red circles.
Figure 6: Ablation study of the trajectory segmentation strategy with different segment numer $N_{s}$ and time theshold $t_{\tau}$. Prompt: a DSLR photo of the Imperial State Crown of England.
@@
Quantitative Comparison. We evaluate the text alignment and human preference of each text-to-3D method as follows: we generate 3D models from 40 prompts selected from the DreamFusion Gallery  [^1] https://dreamfusion3d.github.io/gallery.html, and calculate the average CLIP Score and ImageReward (IR) score across 120 camera views for each model. Following Connect3D [22], we compute the FID score between 40 objects (with 120 rendered views for each) and 50,000 images generated by Stable Diffusion 2.1. As shown in the results, our method achieves the best performance in text alignment, human preference, and fidelity, while requiring the shortest generation time, demonstrating its superior overall performance. The 40 prompts used are provided in the supplementary material.
User Study. We conduct a user study involving 40 volunteers to evaluate the generation quality of each algorithm. Each volunteer is asked to review eight groups of five 360-degree rendered videos and to rank them based on three criteria: text alignment (Q1), object realism (Q2), and object details (Q3; e.g., artifacts, brightness, and the Janus problem). To minimize bias, the order of the videos is randomized. As shown in Tab. 2, users consistently prefer our method across all criteria.
6 Ablation Study
@@
DreamFusion[31] 28.47 -0.004 140.84 60 4.73 4.87 4.90
LucidDreamer[23] 29.99 0.006 121.80 45 2.93 2.98 2.93
Consistent3D[48] 30.60 0.004 113.61 140 4.14 3.88 3.93
ConnectCD [22] 30.73 0.018 112.61 80 1.63 1.80 1.93
SegmentDreamer 30.88 0.020 110.45 38 1.57 1.47 1.33
[…]
Effect of Segment Number $N_{s}$. We visualize the effect of different segment numbers $N_{s}$ in Fig. 6. As observed, a small $N_{s}$, regardless of the segmentation strategy, consistently results in overly smoothed outputs. As $N_{s}$ increases, the generation quality improves. However, when $N_{s}$ becomes too large (e.g. 10), while more details emerge, the overall model representation becomes vague—a phenomenon also noted in LucidDreamer [23]. We find that $N_{s}=5$ performs well in most cases.
Effect of time threshold $t_{\tau}$. As shown in Fig. 6, a large $t_{\tau}$ leads to overly smoothed results for both segmentation strategies. This is because a limited number of sampling steps in the diffusion process fail to effectively preserve the information of the initial point $\mathbf{z}_{0}$, highlighting the necessity of incorporating multi-step deterministic sampling.
Validation of Approximation Strategy. We analyze the impact of replacing $\boldsymbol{G}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ with $\mathbf{z}_{s_{m}}$ in Fig. 7. As shown, $\boldsymbol{G}_{\boldsymbol{\theta}}(\tilde{\mathbf{z}}_{t}^{\boldsymbol{\Phi}},t,\mathbf{y})$ introduces slight overexposure and geometric distortion. In contrast, $\mathbf{z}_{s_{m}}$ not only produces better generation results but also reduces the training time. This suggests a gap between theory and practice, similar to the omission of U-Net Jacobian in DreamFusion [31].